find two consecutive even integers such that the smaller added to three times the larger gives sum of 38
Found 2 solutions by partha_ban, lynnlo:
Answer by partha_ban(41)

Let x = smaller even integer
So, x+2 = next even integer
By condition, x + 3*(x+2) = 38
x + 3x + 6 = 38
4x = 38 - 6 = 32
x = 32/4 = 8
Therefore, the numbers are 8 and 8+2=10

Answer by lynnlo(4176)

THIS IS AN EXAMPLE ONLY SO YOU COULD TRY TO UNDERSTAND YOUR WORK BETTER:
THE SUM OF 46
LET E=THE SMALLEST NUMBER
(E+2)=======THE LARGEST NUMBER
E+3(E+2)=46
E+3E+6=46
4E=46-6
4E=40
E=40/4
E=10=====THE ANSWER TO THE SMALLER NUMBER
10+2=12==THE ANSWER TO THE LARGER NUMBER
PROOF:
10+3*12=46
10+36=46
46=46
JUST REMEMBER THE "2"MEANS AMOUNT OF CONSECUTIVES
"3"MEANS 3TIMES THE LARGER